A recent survey showed that 65% of U. S. employers were likely to require higher employee contributions for health care coverage. Suppose the survey was based on a sample of 500 companies. Find the 95% confidence interval for the proportion of all companies likely to require higher employee contributions for health care coverage.
Answer #1

Solution :

Given that,

\hat p = 65% = 0.65

1 - \hat p = 1 - 0.65 = 0.35

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha / 2 * \sqrt((\hat p * (1 - \hat p)) / n)

= 1.96 * (\sqrt((0.65 * 0.35) / 500)

= 0.042

A 95% confidence interval for population proportion p is ,

\hat p - E < P < \hat p + E

0.65 - 0.042 < p < 0.65 + 0.042

0.608 < p < 0.692

(0.608 , 0.692)